on a grid draw all possible principal angles in standard position,
tanA=1/2

Answer by stanbon(75887)

on a grid draw all possible principal angles in standard position,
tanA=1/2
-------------
The tangent is positive in the 1st and 3rd quadrants.
The tangent always = y/x
----------
In the 1st quadrant plot the point (2,1)
Draw a line segment from the origin to (2,1)
That is the terminal side you want.
-------------------
In the 3rd quadrant plot the point (-2,-1)
Draw a line segment from the origin to (-2,-1)
that is the terminal side you want.
-----------------------------------------
A = 26.57 degrees
or
A = 26.57 + 180 = 206.57 degrees
==================================
Cheers,
Stan H.